OpenStudy (anonymous):

find the derivative of f(x)=sqrt(x^2+2x)

Parth (parthkohli):

\[\rm f(x) = \sqrt{x^2 + 2x}\]Chain rule...\[\rm(2x + 2)\cdot{1 \over 2}\left( x^2 + 2x\right)^{-{1 \over 2}}\]

OpenStudy (anonymous):

but have to keep it in the sqrt format

Parth (parthkohli):

Simplify what I gave you.

OpenStudy (anonymous):

how would i rewrite that in the sqrt format